Run Group C

From clas12-run
Jump to: navigation, search
[edit]

NOTE: All NON-JLab numbers MUST be dialed with (757) area code. When calling from a CH landline, dial "9" first. To call from outside, all 4-digit numbers must be preceded by 757-269

Person Phone Number
Run Coordinator 9-757 575-7540 (cell)
Physics Division Liaison 9-757 876-1789 (cell)
Run Group C coordinator 9-757 639-6640 (cell)
Shift Expert 5244 or 9-757 329-4846 (cell)
Shift Worker 5126
MCC-OPS 7048
Crew Chief 7045
Crew Chief 9-757 876-3367 (cell)
Program Deputy 9-757 876-7997 (cell)
Engineering on-call 9-757 748-5048 (cell)
Engineering on-call secondary 9-757 897-9060 (cell)
Radcon (needed for target cell swaps) 9-757 876-1743 (cell)
Gate House Guard 5822
Location Phone Number
Hall-B Floor 5165
Hall-B Space Frame 5170, 5171
Hall-B Forward Carriage 5371
Hall-B Counting Room 5245 (Shift1), 5126 (Shift2)
Hall-B Counting Room 5244 (Expert)
Hall-B Gas Shed 7115
System Phone Number On-Call Person
DC (757) 748-5048 (cell) Engineering On-Call
SVT + MM (757) 541-7539 (cell) Yuri Gotra Primary
SVT + MM (757) 753-7769 (cell) Rafo Paremuzyan (secondary)
ECAL (757) 810-1489 (cell) Cole Smith
FTOF/CTOF (757) 344-7204 (cell) Jose Carvajal
CND (757) 329-4844 (cell) Pierre Chatagnon
HTCC-LTCC (757) 344-7174 (cell) Youri Sharabian
FT (757) 344-1848 (cell) Raffaella De Vita
BAND (757) 310-7198 (cell) Florian Hauenstein
RICH (757) 344-3235 (cell) Valery Kubarovsky
DAQ (757) 232-6221 (cell) Sergey Boiarinov
Slow Controls (757) 748-6922 (cell) Nathan Baltzell
Beamline (757) 303-3996 (cell) Eugene Pasyuk
Pol. Target On-Call (757)-218-2266 (cell) See schedule or call RC
Polarized Target (757) 746-9277 (cell) Chris Keith
Polarized Target (757) 897-1415 (cell) James Maxwell
Polarized Target (757) 871-5374 (cell) James Brock
Target (not used) (757) 822-9586 (cell) Bob Miller Primary
Target (not used) (757) 897-9060 (cell) Denny Insley Secondary
Counting House C x6000
Counting House D x5504

In case of a problem with superconducting magnets call Engineering on-call first, particularly if it is related to cryo.
For all other superconducting magnet problems call magnet group.
Their schedule and phone numbers are posted on the wall next to the Saclay target computer.


List of people authorized to operate the SC magnets in Hall-B

  • Hall B - Denny Insley, Morgan Cook, and Eugene Pasyuk
  • Target Grp - Chis Keith and James Maxwell
  • Magnet Grp - Probir Ghosha and Renuka Rajput-Ghoshal
  • DSG - Brian Eng, Pablo Campero and Tyler Lemon
  • DC Power - Onish Kumar and Sarin Philip




Click Here to edit Phone Numbers. Note, you then also have to edit the current page to force a refresh.


CLAS12 Run Group C

Run Coordinator: Stuart Fegan (9-)757 575-7540

PDL: Eugene Pasyuk (9-)757 876-1789 , x6020 office

Run Group Coordinator: Sebastian Kuhn (9-)757 639-6640

Daily RC Meetings at 14:00: CH Room 200C, and on Zoom Zoom information: Link, Meeting ID: 160 867 4253, Passcode: 786506

Run Coordinator and Target Expert Schedule

Longterm Run Plan

Medium-Term Run Plan (FTon Portion Part 2)

IMPORTANT SHIFT NOTES

* In case of medical or fire emergency, call 9-911 from a LANDLINE phone in the Counting House
* If you have symptoms or a positive COVID test, do NOT come to your shift - instead, inform the PDL and the RC and JLab medical services immediately so that replacements can be arranged.
* Remote shift: If needed, worker shifts can be taken from off-site. Inform the PDL and RC. Also, make sure that you follow the instructions under the "Worker Shift" Tab - you need some lead time to set up!
Note to shifts, please:
* read the logbook, be aware of previous 24 hours
* arrive at least 10 minutes early for shift change
* refresh the run wiki pages, look over them every shift:
** follow the expert/worker shift instructions 
** find latest pitfalls and known issues under GENERAL INFORMATION
* communicate with your co-shift person
** the shift expert should know about and refer to logbook entries in their shift summary
** the shift worker should discuss an issue with the shift expert before making extra logbook entries

Short Term Schedule for the Week from Wednesday, February 1 - 8, 2023

  • Data taking in FTon configuration, inbending torus field, 5-pass beam on ND3.
  • Current: 4 nA
  • Run duration: 100M events or 4 hours, whichever comes first
  • Note, BOM is not installed for this experiment, so it does not matter whether it appears masked in FSD or not.
  • New conditions for raster operation:
    • IF there is any ammonia sample (NH3 and ND3) in the target, the raster MUST be operating at all times. The Raster FSD must be UNMASKED. Orbit locks MUST be OFF (except for brief moments to restore beam position).
    • For beam tuning or other non-rastered beam, the Raster FSD must be MASKED.
    • If MCC needs to diagnose beam with a harp, we can offer them tagger 2C24 and follow the procedures required for that. 2H01 harp scan should not be done when we have beam raster on.
    • Changing raster offset: ask MCC to NOT correct whatever happens on 2H01. Note new 2H01 values and inform MCC of the new 2H01 x and y values - then they can keep the beam at those positions.
  • Experiment setup is "FTon"
  • If HWP is changed, start a new run

PLAN for Wednesday, February 1, 2023

Evening Shift

  • Restore beam on tagger initially
  • Take production data with 4 nA on positively polarized ND3 (note the HWP state)

PLAN for Thursday, February 2, 2023

Night Shift

  • Take production data with 4 nA on positively polarized ND3

Day Shift

  • Take production data with 4 nA on positively polarized ND3

Evening Shift

  • Take production data with 4 nA on positively polarized ND3

PLAN for Friday, February 3, 2023

Night Shift

  • Take production data with 4 nA on positively polarized ND3

Day Shift

  • HWP flip at start of day shift (after MCC meeting)
  • Take production data with 4 nA on positively polarized ND3

Evening Shift

  • Take production data with 4 nA on positively polarized ND3

PLAN for Saturday, February 4, and Sunday, February 5, 2023

  • Take production data with 4 nA on positively polarized ND3
  • Switch to solid target Sunday at 1pm
  • Production running with solid targets through Tuesday

Sunday Swing Shift

  • Production running on CD2 at 4 nA

PLAN for Monday, February 6, 2023

Night Shift

  • Take production data on CD2 at 4 nA

Day Shift

  • Take production data on CD2 at 4 nA


NOTES: 
* Watch the beam positions (and thus the midstream halo counters) closely and ask MCC to adjust the positions if they are (on average) >=0.1mm off.
* Remember, the hall will only be dropped in controlled mode once someone goes downstairs and enters the hall. You/RadCon should do that right away after calling MCC. 

DAQ configuration

  • Configuration is currently PROD66/PROD66_PIN, trigger file rgc_300MeV_v1.4_no_DC.trg
  • NOTE: UNLESS a RICH problem crashes the DAQ, do NOT interrupt an otherwise smooth run to do a RICH recovery - instead do this in between runs.
  • NOTE: BOTH the shift expert AND the shift worker MUST check and reset monitoring histograms at least once every 30-60 minutes to make sure there are no major detector problems. Sometimes it is the only way to know that a roc needs to be rebooted. Watch the raster pattern!
  • At the end of each run, follow the STANDARD DAQ RESTART SEQUENCE:
    • "end run".
    • If the run ended correctly and you will run with the same configuration and trigger file, then: "prestart", "go"
    • If you are changing configuration (between PROD6X and PROD6X_PIN) or trigger file: "cancel", "reset", "configure", "download", "prestart", "go".
    • if the run did not end correctly or if any ROCs had to be rebooted:
      • Quit ALL running instances of mon12 and rastermon
      • "cancel", "reset"
      • "configure", "download", "prestart", "go"
      • restart mon12 and rastermon
  • After each step, make sure it is completed in the Run Control message window. If a ROC has crashed, find which one it is, issue a roc_reboot command ON JUST THAT ROC and try again. Contact the DAQ expert if there are any questions.

References and Standards

Values will need to be updated (Jan. 30, 2023)

5-pass (10.55 GeV) Running conditions

  • Torus at 100% current, inbending electrons. Setpoint = +3770 A (Readback = +3766 A)

Solenoid set-point of 2417.5 A results in the read back 2420.0 A and exactly 5.00 T field on center


Raster settings (current as of February 2, 2023)

  • x: Vpp = 2.75 V, offset = 0.0 V
  • y: Vpp = 2.05 V, offset = 0.28 V
  • Fundamental Period = 2 s
  • The correct raster calibration constants should be automatically loaded by rastermon - if not, enter them through "Config" by hand:
  • f(x) = 0.0104*x + -25.5
  • f(y) = -0.0115*y + 24.4

Nominal Beam Positions (current as of February 2, 2023)

  • 2C21A, X = -0.4 mm, Y = +0.6 mm
  • 2C24A, X = -2.0 mm, Y = +0.6 mm
  • 2H01, X = -1.0 mm, Y = +0.19 mm
  • Orbit lock on 2C24 and 2H01 ONLY if raster is OFF, otherwise orbit lock on 2C21 and 2C24 (note: this appears as "stopped" on the beam line GUI).

FSD Thresholds

  • Upstream: 2000 Hz
  • Midstream: 2000 Hz
  • Downstream: 100000 Hz (1e5)
  • BOM: 1000000 Hz (BOM not installed for RG-C)
  • 5 ms dwell time

Reference plots and histograms

Counter rates

  • With 6 nA beam, the upstream counters should read zero.
  • The Midstream ones should read 0-20 Hz or so.
  • Greatly increased rates may indicate bad beam tune or bleed-through from other Halls.

see below for Known Detector Issues

General Information

  • Controlled access: after you call MCC, they will drop the hall to controlled access once someone goes down and calls them for entry. Then call when they should go back to beam permit and call again to say you want the beam back and what beam current. Rapid access is controlled access without RadCon.
  • If MCC needs to diagnose beam with a harp, we can offer them tagger 2C24 and follow the procedures required for that. 2H01 harp scan should not be done when we have beam raster on.
  • BMT:
    • After HV trip: the channel HV should be turned back on as soon as you notice the trip, power cycle by turning HV OFF and ON without waiting for any amount of time.
    • HV frequently tripping: Call expert. Do not change HVs without approval from the expert.
  • mon12: It can happen that histograms do not get cleared completely at run start. This is seen eg as an excess of small multiplicities. The remedy is to quit mon12 and open it again. Other than that, small multiplicities come from time intervals with beam down. You might want to look at the multiplicities before uploading the mon12 histograms in the logbook.
  • DAQ:
    • clondaq6/data is full, switched to clondaq7, rastermon should be using ET on clondaq7
    • It should no longer be necessary to quit RasterMon at the end of each run and restart after a successful "Go" for the next run. However, if DAQ crashes, you SHOULD follow these steps, and the same for mon12.
  • Magnet trip: call "Engineering on call", if due to cryo, also inform the target expert about the cryogenic issues.
  • CND CAMAC crate was off? Need to run the script to set the thresholds for CND CFDs.
  • Hall lights: Before leaving the hall and going to beam permit check that dome lights, forward carriage lights, and pie tower lights are OFF.
  • Moller runs: should be taken with 8-10 nA.
  • Beam viewer: Please REMOVE the beam viewer unless we are actively tuning the beam or calibrating beam position/raster, to avoid damage to electronics, cameras etc.
  • Change of running conditions
    • Check that the trigger alarms (on rate/nA) are set appropriately and RC changes them as needed.
    • Whenever the target type or polarization direction changes, or the half-wave plate status changes, make sure to enter a comment in the BEGIN RUN / END RUN comment box.
    • Do not change running conditions while a run is ongoing. This includes changing target polarization or HWP, etc.

Known Issues

Fr 11 Nov 2022 ECAL PMT_SEC2_WO8 HV is disabled https://logbooks.jlab.org/entry/4080950

Follow-up Re: RICH tile 21 in sector 1 is dead. Working without one tile out of 276 for a moment https://logbooks.jlab.org/entry/4079453

Run #17396: RICH lost three scalars https://logbooks.jlab.org/entry/4079190

tbc

Shift expert tasks

  • Come to your shift in time for the shift change, 10-15 minutes before the start of your shift.
  • Follow the Short Term Schedule as outlined by RC
    • Run Plan, DAQ Configuration, and References and Standards
    • also check tab GENERAL INFORMATION
  • Keep shift summary up to date in HBLOG. Record all that happens.
  • Polarized target:
    • Helicity-Latched Trigger GUI and the Helicity Trigger Asymmetry strip-chart (not accessible remotely) (example https://logbooks.jlab.org/entry/4041939). (ONLY WHILE ACTUALLY RUNNING ON NH3 AND ND3 targets!)
    • In case of any issues, contact the polarized target expert on call.
    • Ice on the run valve
    • NMR running (ONLY WHILE ACTUALLY RUNNING ON NH3 AND ND3 targets!)
      • The JLab Polarization Display is seen via a VNC window in the counting room. Here, the NMR Controls' blue bar must be moving, indicating that the NMR is running and the polarization is measured.
      • The NMR frequency needs to be controlled by hand. If you see target polarization (on the red graph) starting to drop, call the expert. They will guide you through the adjustment procedure. This will be in the electronics room by the door to the counting room, in the third crate from the right ("CR-6"). White box which says "EIO Tube" and has a red button at the top right, which says "Move", next to a display for "Position". This will be the one to press, under guidance of the target expert.
    • Alarms from System RG-C Polarized Target (alarm handler).
  • Do a RICH recovery (press the RICH Recovery button on the GUI) once daily during day shift, between runs. Please log in shift summary.
    • Note the automated "RICH Recovery" sequence already confirms RICH4 is up and accepting ssh connections before claiming success.
    • You have to "Cancel->Reset->Configure->Download->Prestart" before starting the next run.
  • Fill out BTA hourly. Click "Load from EPICS" to automatically fill the left side then correct as needed.
    • Accelerator beam time:
      • ABU: acceptable beam used (eg harp scans, calibrations, trigger settings, running)
        • The info from EPICS is not always correct, e.g. when taking data with low beam current, you should change the ABU to the proper (estimated) time by hand instead of simply accepting "0" if you are actually taking data.
      • BANU: beam acceptable but not used (eg radiator change or our problems)
      • BNA: beam not acceptable (eg tuning, drifting, not focused)
    • Experiment beam time
      • ER: experiment ready
      • PCC: planned configuration change (when beam is masked while radiators are moved, DAQ (stopping/starting runs, tests, ...) are PCC..
      • UED: unplanned experiment down, is used for hardware failure
    • Enter fields at tab "Shift Information"
    • Promptly send "Info for MCC" at 7:00, 15:00, and 23:00
    • Sign, viz log completed 8 hour sheet


In case of other questions or uncertainties of what to do, call the RC or PDL

(remote) shift worker: see dedicated Worker Shift tab above.


Shift Worker Tasks

  • Check the information in the Short Term Schedule and GENERAL INFORMATION, for every shift
  • Come to your shift in time for the shift change, 10-15 minutes before the start of your shift.
  • Use the new HBRUN logbook for automated and run-specific entries and screenshots, see https://logbooks.jlab.org/entry/4048698


Make the following log entries ONCE per run

  • Screenshots: submit to logbook only if you capture the information with beam ON
    • Beamline Overview, in VNC or Web Browser
    • CLAS12 Trigger rates and DAQ status
    • RGC Polarized Target GUI - under Beamline on CS-Studio
    • Helicity Asymmetry Strip Charts - see [11] for an example
  • Upload monitoring plots to logbook
    • Standard spectra (mon12): inspect all monitoring histograms ' and carefully compare them to the corresponding reference histograms: reference, current, and 5 most recent runs, see link under tab Useful Links
      • You can choose a different reference run if you are confident what you are doing, the reference numbers refer to logbook entries.
    • Raster Monitoring (use "Log Entry" button on RasterMon screen at bottom-right corner).
      • An example of a "decent looking" RasterMon is here: [12]. The x and y distributions should show a smooth dome, without any horns or spikes on it and without the distributions being cut off at the edges.

Once per shift:

  • Fill the shift checklist
  • Log screen shots of RICH temperature and scaler maps. Log in RICH logbook.

Regularly:

  • reset & check spectra frequently (every 30 - 60 minutes)
    • Some plots in the online mon12 have a choice of histogram sets per sector, select buttons at the bottom of the plot. That means you have to look at all sectors for these histograms.
    • The logbook entry shows all histograms.
  • Check the life cam of the polarized target run valve for plumes or ice buildup (hallbcam04)
    • Make sure you refresh the web page! Its no longer necessary to put a snapshot into the logbook UNLESS something looks suspicious. In THAT case, immediately inform the target expert(s).
  • Log the strip charts in time intervals according to the x-axis time span, via right-click on the screen and selecting "Make Elog Entry". Make sure NOT to submit to ELOG.
    • Beam current and position strip-charts (CS-Studio -> Beamline -> Strip Charts)
    • Helicity Trigger Asymmetry strip-chart (CS-Studio -> Helicity-Latched Trigger GUI Screen -> Menu -> myaPlot) (not accessible remotely) (ONLY WHILE ACTUALLY RUNNING ON NH3 AND ND3 targets!
      • Example
      • Choose reasonable y / x axis ranges: R-click on the axis -> Modify (axis-name) Axis
    • For the y axis in the bottom panel, choose the target polarization and a reasonable range for the y axis
      • Set the time axis auto-update. A ">>" symbol will appear below the time axis when this is activated.
        • Right-click on X axis -> Set Auto Time Stepping

For REMOTE Shifts (offsite):

ATTENTION: If you plan to take (a) remote shift(s), you MUST test your setup (including VNC connection) at least 24 hours before your first shift!


Clas12design.png